# Cellular, Cell-less, and Everything in Between: A Unified Framework for Utility Region Analysis in Wireless Networks

### v2

@@ VI Summary
In modern affine interference models that take self-interference into account, we have demonstrated that the SINR and rate regions are convex if the interference matrices are inverse Z-matrices. Moreover, we have established numerous connections between interference patterns in wireless networks and the extensive mathematical literature on these matrices to facilitate the identification of convex SINR and rate regions. Furthermore, we disproved a conjecture posed in [13, Sect. 5.4.4][24], and we introduced a new concept of channel Z-compatibility for users that addresses certain limitations of previous notions of favorable propagation in the massive MIMO literature. Our findings also provide both numerical and theoretical support for formulating sum-rate maximization problems using rates, rather than SINR levels, as the optimization variables. By doing so, we increase the chance of uncovering hidden convexity in the problems.
While the aforementioned results involving inverse Z-matrices require interference models where beamformers remain unchanged if the transmit power changes, much of our developed theory also extends to scenarios where this assumption does not hold. In particular, we have derived simple characterizations of the SINR and rate regions using the concept of the spectral radius of nonlinear mappings. This approach simplifies notation and unlocks powerful machinery with a long history in mathematics, potentially enabling further extensions of the current results in future studies.
@@
[30, Theorem 2.9.2] A nonnegative matrix ${\bm{M}}\in{\mathbb{R}}^{N\times N}_{+}$ is an inverse Z-matrix for $N=2$ if and only if $\mathrm{det}({\bm{M}})>0$.
@@
[30, Theorems 4.5 and 9.2] Let ${\bm{M}}\in{\mathbb{R}}_{++}^{N\times N}$. Then there exists $\alpha>0$ such that $\alpha{\bm{I}}+{\bm{M}}$ is an inverse Z-matrix.
@@
[30, Theorem 1.7] If ${\bm{M}}\in{\mathbb{R}}^{N\times N}$ is an inverse M-matrix, then ${\bm{M}}+{\bm{D}}$ is also an inverse M-matrix for any nonnegative diagonal matrix ${\bm{D}}\in{\mathbb{R}}_{+}^{N\times N}$.
@@
[30, Theorem 2.1] Let the nonnegative matrix ${\bm{M}}\in{\mathbb{R}}_{+}^{N\times N}$ be invertible. Then $\bm{M}$} is an inverse Z-matrix if and only if ${\bm{M}}^{-1}$ is an M-matrix (i.e., $\bm{M}$} is an inverse M-matrix).
@@
[28, Theorem 4.3] Assume that the nonnegative matrix ${\bm{M}}\in{\mathbb{R}}_{+}^{N\times N}$ is an inverse Z-matrix. Then the mapping $\mathcal{D}_{+}^{N\times N}\to{\mathbb{R}}_{+}:{\bm{D}}\mapsto\rho({\bm{DM}})$ is a convex function.
@@
The next result is immediate from [29, Lemmas 3.2 and 3.3] and the definition of the nonlinear spectral radius.
@@
[29, Theorem 3.1(2)] Fix a norm $\|\cdot\|$ in ${\mathbb{R}}^{N}$, and define $\mathcal{P}:=\{{\bm{x}}\in{\mathbb{R}}_{+}^{N}\mid\|{\bm{x}}\|=1\}$. For a sequence $(G_{n}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}^{N})_{n\in{\mathbb{N}}}$ of general interference mappings, assume that there exists a general interference mapping $G:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}^{N}$ such that $\lim_{n\to\infty}\sup_{{\bm{x}}\in\mathcal{P}}\|G_{n}({\bm{x}})-G({\bm{x}})\|=0$. Then $\lim_{n\to\infty}\rho(G_{n})=\rho(G)$.
@@
[23, Proposition 4] Let $T:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{++}^{N}$ be a standard interference mapping and $T_{\infty}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}^{N}$ be its asymptotic mapping (see Definition 2). Then $\mathrm{Fix}(T)\neq\emptyset$ if and only if the (nonlinear) spectral radius $\rho(T_{\infty})$ of $T_{\infty}$ satisfies $\rho(T_{\infty})<1$.
@@
[23, Proposition 5] Let $T:{\mathbb{R}}^{N}_{+}\to{\mathbb{R}}_{++}^{N}$ be a standard interference mapping and $\|\cdot\|$ a monotone norm. Then $T$ has a fixed point $\bm{x}$$^{\star}$ satisfying $\|{\bm{x}}^{\star}\|\leq 1$ if and only if the conditional eigenvalue $\lambda$ associated with $T$ and $\|\cdot\|$ satisfies $\lambda\leq 1$.
Proof of Lemma 1
Proof.
Proof of the statement in Example 1
Proof.
Standard Perron-Frobenius theory shows that the spectral radius of $\rho(G)$ is positive, and there exists a positive left eigenvector ${\bm{l}}\in{\mathbb{R}}^{N}_{++}$ and a positive right eigenvector ${\bm{r}}\in{\mathbb{R}}_{++}^{N}$ associated with $\rho(G)$. Choose $\epsilon>0$ such that ${\bm{M}}_{1}={\bm{M}}-\epsilon~{\bm{lr}}^{t}\in{\mathbb{R}}_{+}^{N\times N}$ (a scalar $\epsilon$ with this property always exists because $\bm{M}$} is positive). The mapping $T:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{++}^{N}:{\bm{x}}\mapsto{\bm{M}}_{1}{\bm{x}}+\epsilon{\bm{l}}$ is a standard interference mapping because it is positive and concave [39, Proposition 1], and it satisfies hypothesis ($T{T_{\|\cdot\|}}$), by construction, for any monotone norm $\|\cdot\|$. In particular, using the monotone norm $(\forall{\bm{x}}\in{\mathbb{R}}^{N})~\|{\bm{x}}\|_{\bm{r}}:={\bm{r}}^{t}|{\bm{x}}|$, we have $(\forall{\bm{x}}\in{\mathbb{R}}_{+}^{N})~{T_{\|\cdot\|}}({\bm{x}})=G({\bm{x}})={\bm{Mx}}={\bm{M}}_{1}{\bm{x}}+\epsilon{\bm{l}}\|{\bm{x}}\|_{\bm{r}}$. Therefore, we are in the setting of Definition 5 because $g_{G}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}^{N}:{\bm{x}}\mapsto\rho(\mathrm{diag}({\bm{x}})G)=\rho(\mathrm{diag}({\bm{x}}){\bm{M}})$ is convex on ${\mathbb{R}}_{+}^{N}$ as a direct consequence of Fact 5 in the Supplemental Material, so $G$ is norm inducing as claimed, and the induced norm satisfies $(\forall{\bm{x}}\in{\mathbb{R}}_{+}^{N})\|{\bm{x}}\|_{G}=\rho({\mathrm{diag}(\bm{x})}{\bm{M}})$. ∎
Proof of the statement in Example 2
Proof.
[…]
Proof of Proposition 6
Proof.
Now, follow the exact same steps below (25) to conclude that the set $\overline{\mathcal{R}_{\mathcal{P}}}$ is convex. Boundness of ${\bm{r}}^{\star}\in\overline{\mathcal{R}_{\mathcal{P}}}$ follows immediately from boundness of closed balls $\{{\bm{x}}\in{\mathbb{R}}^{N}\mid\|{\bm{x}}\|_{T_{\|\cdot\|}}\leq 1\}$, nonnegativity of vectors in $\overline{\mathcal{R}_{\mathcal{P}}}$, and the fact that $x\to\infty$ implies $e^{x}\to\infty$. Therefore, being closed and bounded in the standard (finite-dimensional) Euclidean space, $\overline{\mathcal{R}_{\mathcal{P}}}$ is compact in this space. ∎

### v3

@@ VI Summary
We have derived simple characterizations of the SINR and rate regions, with and without power constraints, using the concept of spectral radius of nonlinear mappings. This viewpoint streamlines the analysis and connects the study of feasible regions in wireless networks to a well-developed mathematical framework, opening the door to further extensions. In particular, our characterization of the feasible rate region as the level-set of a function involving the nonlinear spectral radius enables the identification of regimes in which time sharing brings no benefit with respect to the achievable rates. In particular, for affine interference models that explicitly account for self-interference, we show that convexity of the SINR and rate regions is guaranteed under a tractable sufficient condition: the relevant interference matrices are inverse Z-matrices. We further connect common interference patterns in wireless networks to the extensive theory of inverse Z-matrices, providing tools to recognize convex feasible regions without explicitly computing matrix inverses. Beyond these structural results, we disprove the conjecture posed in [35, Sect. 5.4.4][36], and we introduce the notion of channel Z-compatibility, which overcomes some limitations of the concept of favorable propagation in the massive MIMO literature. Finally, supported by both theory and numerical evidence, we advocate formulating sum-rate maximization problems in terms of rates rather than SINR levels, as this choice can reveal hidden convexity and lead to tractable large-scale optimization problems.
@@
[19, Theorem 2.9.2] A nonnegative matrix ${\bm{M}}\in{\mathbb{R}}^{N\times N}_{+}$ is an inverse Z-matrix for $N=2$ if and only if $\mathrm{det}({\bm{M}})>0$.
@@
[19, Theorems 4.5 and 9.2] Let ${\bm{M}}\in{\mathbb{R}}_{++}^{N\times N}$. Then there exists $\alpha>0$ such that $\alpha{\bm{I}}+{\bm{M}}$ is an inverse Z-matrix.
@@
[19, Theorem 1.7] If ${\bm{M}}\in{\mathbb{R}}^{N\times N}$ is an inverse M-matrix, then ${\bm{M}}+{\bm{D}}$ is also an inverse M-matrix for any nonnegative diagonal matrix ${\bm{D}}\in{\mathbb{R}}_{+}^{N\times N}$.
@@
[19, Theorem 1.1] Let the nonnegative matrix ${\bm{M}}\in{\mathbb{R}}_{+}^{N\times N}$ be invertible. Then $\bm{M}$} is an inverse Z-matrix if and only if ${\bm{M}}^{-1}$ is an M-matrix (i.e., $\bm{M}$} is an inverse M-matrix).
@@
[17, Theorem 4.3] Assume that the nonnegative matrix ${\bm{M}}\in{\mathbb{R}}_{+}^{N\times N}$ is an inverse Z-matrix. Then the mapping $\mathcal{D}_{+}^{N\times N}\to{\mathbb{R}}_{+}:{\bm{D}}\mapsto\rho({\bm{DM}})$ is a convex function.
@@
The next result is immediate from [28, Lemmas 3.2 and 3.3] and the definition of the nonlinear spectral radius.
@@
[28, Theorem 3.1(2)] Fix a norm $\|\cdot\|$ in ${\mathbb{R}}^{N}$, and define $\mathcal{P}:=\{{\bm{x}}\in{\mathbb{R}}_{+}^{N}\mid\|{\bm{x}}\|=1\}$. For a sequence $(G_{n}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}^{N})_{n\in{\mathbb{N}}}$ of general interference mappings, assume that there exists a general interference mapping $G:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}^{N}$ such that $\lim_{n\to\infty}\sup_{{\bm{x}}\in\mathcal{P}}\|G_{n}({\bm{x}})-G({\bm{x}})\|=0$. Then $\lim_{n\to\infty}\rho(G_{n})=\rho(G)$.
@@
[9, Proposition 4] Let $T:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{++}^{N}$ be a standard interference mapping and $T_{\infty}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}^{N}$ be its asymptotic mapping (see Definition 2). Then $\mathrm{Fix}(T)\neq\emptyset$ if and only if the (nonlinear) spectral radius $\rho(T_{\infty})$ of $T_{\infty}$ satisfies $\rho(T_{\infty})<1$.
@@
[9, Proposition 5] Let $T:{\mathbb{R}}^{N}_{+}\to{\mathbb{R}}_{++}^{N}$ be a standard interference mapping and $\|\cdot\|$ a monotone norm. Then $T$ has a fixed point $\bm{x}$$^{\star}$ satisfying $\|{\bm{x}}^{\star}\|\leq 1$ if and only if the conditional eigenvalue $\lambda$ associated with $T$ and $\|\cdot\|$ satisfies $\lambda\leq 1$.